OpenStudy (liv33):

What is the solution of m/4 - 3 = 7 ?

Mehek (mehek14):

Add 3 m/4 = 7 + 3

OpenStudy (liv33):

10

Mehek (mehek14):

Multiply by 4 m = 10 * 4

OpenStudy (liv33):

40

Mehek (mehek14):

m = 40

OpenStudy (liv33):

Thanks again!

Mehek (mehek14):

You're welcome :) Are you understanding this?

OpenStudy (liv33):

like 80%

Mehek (mehek14):

Ok (:

Mehek (mehek14):

I'll try to get 100% cx